THE ROTATIONAL SPECTRUM AND OBSERVATIONAL STUDY OF HYDROXYACETONE

\underline{ROGIER BRAAKMAN}, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125; BRIAN J. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; SUSANNA L. WIDICUS, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125; and GEOF-FREY A. BLAKE, Division of Geological and Planetary Science, California Institute of Technology, Pasadena, CA 91125.

The rotational spectrum of hydroxyacetone, or acetol, has been studied up to 346 GHz. Due to internal rotation of the methyl top, the spectrum is split into A and E states. For the A state, which was treated as a semi-rigid molecule, 578 new lines have been assigned and fit in the ground state up to J=60. Precise rotational and centrifugal distortion constants have been determined for this state. Using the splitting between the A and E states as a guide, 288 new lines in the ground E state have been assigned up to J=60 as well. Fitting of these lines is currently underway. Because of its close structural relationship to glycolaldehyde and dihydroxyacetone, both of which have been detected in the galactic center region SgrB2(N-LMH)^{a}^{b}, an observational search of acetol was conducted in this same region at the Caltech Submillimeter Observatory in july 2004. No lines were successfully observed, but a column density upper limit of 8.0×10^{14} cm^{−2} was determined.

aJ. M. Hollis, F. J. Lovas and P. R. Jewell, 2000, ApJ 540, 107-110.

bS. L. Widicus-Weaver and G. A. Blake, 2005, ApJ, submitted.